OpenStudy (anonymous):

D = {[x, f(x)]: (-1, 3), (0, 2), (1, 3), (2, 6), (3, 11)} Write the rule for f(x).

OpenStudy (zzr0ck3r):

hint (-1)^2+3 = 2 0^2+2 = 2 3^2+2 = 11

OpenStudy (anonymous):

f(x) = x^2 +2